OpenStudy (anonymous):

evaluate 8 − 12 ÷ 2 + 3

OpenStudy (anonymous):

-4/5

OpenStudy (ivettef365):

no first do the division, PEMDAS

OpenStudy (irishboy123):

do division first - BIDMAS/BODMAS becomes 8-6+3 simples

OpenStudy (anonymous):

\[8 - \frac{12}{2} + 3 = 8-6+3 = ...\]

OpenStudy (nirmalnema):

1

OpenStudy (anonymous):

I don't think your answer is right @nirmalnema

OpenStudy (ivettef365):

what is 11 - 6 = ??

OpenStudy (nirmalnema):

its 5

OpenStudy (ivettef365):

good